OpenStudy (anonymous):

how many points of intersections are there between the two functions? √x and 3|x|+1

OpenStudy (michele_laino):

please note that: \[y=3|x|+1=3x+1, x \ge 0\] and: \[y=3|x|+1=-3x+1, x<0\]

OpenStudy (anonymous):

thanks

OpenStudy (michele_laino):

now, in order to find intersection between y=3x+1 and y=sqrt(x), you have to solve this equation: \[3x+1=\sqrt{x}\] whereas if you want to find intersection between -3x+1 and sqrt(x), you have to solve this equation: \[-3x+1=\sqrt{x}\]